[figure: 9.7 cm]

MAGNETIC MOMENT OF Fe IN RFe_{11}Ti (R = Y, Ho, Dy) INTERMETALLICS UNDER HIGH PRESSURE

O. Mikulina^{a}, J. Kamarád^{a}, Z. Arnold^{a}, C. Abadia^{b}, M.R. Ibarra^{b}, P.A. Algarabel^{b}

aInstitute of Physics AS CR, Cukrovarnická 10, 162 53 Praha 6, Czech Republic

bDpto. de Fisica de la Materia Condensada – ICMA, Facultad de Ciencias, Universidad de Zaragoza – CSIC, 50009 Zaragoza, Spain

The Fe–rich rare earth intermetallics exhibit extraordinary magnetovolume effects that are ascribed to volume instabilities of the Fe–moment m_{Fe}. Recently, a theoretical analysis elucidated the effect of the local Fe–environment on m_{Fe} in the family of Y–Fe intermetallics [1]. Conclusions of this analysis have been used to interpret the different pressure dependencies of m_{Fe} observed in the Fe–rich YFe_{11}Ti and Y_{2}Fe_{17} intermetallics [2]. We report the magnetization measurements on RFe_{11}Ti (R = Y, Ho, Dy) under pressure up to 1 GPa and at fields up to 5 T performed using a CuBe pressure cell in a SQUID magnetometer. The decrease of m_{Fe} under pressure depends very slightly on the R–elements in RFe_{11}Ti, where we have obtained d ln m_{Fe}/dp|_{5K} = −0.030 GPa^{−1} and −0.034 GPa^{−1} in RFe_{11}Ti (R = Ho, Dy) and YFe_{11}Ti, respectively. This result contrasts with the strong dependence of d ln m_{Fe}/dp|_{5K} on the R–elements in the R_{2}Fe_{17} intermetallics [3].

[1] R.F. Sabiryanov, S.S. Jaswal, Phys. Rev. B 57, 7767 (1998)

[2] O. Mikulina et al., J. Magn. Magn. Mats. - in press (1999), {EMMA 98, Zaragoza} [3] J. Kamarád et al., J. Appl. Phys. - in press (1999), {MMM 98, Miami}

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Dr. Jiri Kamarád

Address for correspondence : Institute of Physics, AS CR Cukrovarnická 10 162 53 Praha 6 Czech Republic

Email address : kamarad@fzu.cz